# User Simulation for Evaluating Information Access Systems

### v1

@@ 2.3 User Modeling
User simulation can be regarded as developing a complete and operational user model, which makes work on user modeling highly relevant to us.
@@
Within the broader field of user modeling, the most relevant lines of work for include the following: (1) Search tasks and intent, including, e.g., typology of search tasks [Marchionini:2006:ACM], taxonomy of Web search [Broder:2002:SIGIRForum], and taxonomy of e-commerce queries [Sondhi:2018:SIGIR]. These studies provide a basis for studying simulation of different kinds of search tasks and intent. (2) Information seeking models, including, e.g., exploratory information seeking models [Ellis:1989:JD, Meho:2003:JASIST]. Such models provide guidance on high-level workflow of a simulated user when performing an information seeking task. (3) Cognitive models of users, which attempt to model the knowledge state of a user; for example, the ASK theory [Belkin:1982:JD] and Ingwersen:1996:JD’s cognitive IR model [Ingwersen:1996:JD]. (4) Contextual user modeling[Ingwersen:2005:book], which takes a broader view of the context that may affect a user’s information seeking behaviour. (5) Click models[Chuklin:2015:Book]that can be directly used in a user simulator to simulate a user’s clicking behaviour. (6) Economic IR models[Azzopardi:2011:SIGIR, Azzopardi:2014:SIGIR], which provide guidance on how users make decisions during an interactive retrieval process and may also be directly used in a user simulator.
As user simulation is essentially to define a computational model for a user, the discussion of user modeling will appear throughout the book. However, as a separate research topic, user modeling clearly plays an increasingly important role as we attempt to optimize the collaboration of AI systems and human users since the objective function to be optimized by an AI system in such an assistive AI application scenario must include formal descriptions of the user with whom the AI system collaborates. As will be elaborated later in the book, so far, the user simulation techniques only use a limited portion of research results on user modeling, leaving much room for future work on incorporating more user modeling research results into user simulation.
@@
In this chapter, we provide a high-level broad overview of research on user simulation with a discussion of the different motivations and applications of how user simulation has been employed in the past. Specifically, we identify three different purposes of studying user simulation: (1) analysis of user behaviour in specific applications, (2) automatic evaluation of systems, (3) synthetic data generation for training models (e.g., using reinforcement learning). While our main interest in this book lies in (2), it is worth pointing out that the first two categories share a lot of similarities in terms of modeling user behaviour in an interpretable way; in contrast, the interpretability aspect may be less critical for third category of application. Moreover, as the research in the first two categories has historically been done mostly in the context of search engine applications, our discussion in the rest of this chapter will also be mostly about search engines, though we will also include a brief discussion of recommender systems and conversational assistants, which have attracted much attention recently. Specifically, in the rest of the chapter, we will first discuss the different motivations for studying user simulation, roughly corresponding to different benefits of user simulation. We then provide an overview of the research done so far in the context of three representative information access systems, i.e., search engines, recommender systems, and conversational assistants. We end the chapter with a discussion of the different applications of user simulation, including particularly the use of simulation for evaluation, which is the main focus of this book.
@@ 3.1 Motivations for User Simulation
First, the study of user simulation has often been done in the context of studying user behaviour, where user simulation can provide multiple benefits, such as formally defining a hypothesis about a user’s behaviour, modeling variations of user factors and their influences on user behaviour, and counterfactual analysis of impact of system components on user behaviour. In general, studying how various factors influence user interaction is important for both understanding users and optimizing the design of user interfaces and algorithms to serve users. For this reason, the study of user behaviour in interactive systems has been the subject of decades of user-oriented IR research (cf. Section 2.3). However, a major challenge in user studies is that the experimental conditions can introduce variables that are difficult to control (e.g., users’ background, domain knowledge, or perseverance), but they can affect how users interact with an information access system. Such uncontrolled variables can be explicitly modeled in user simulations to enable more accurate modeling of a user’s real behaviour.
@@
Moreover, simulation also allows for examining the effects of changes made to various system components on anticipated user behaviour. It facilitates a more robust assessment than what a live environment would typically allow (e.g., estimating long-term engagement/satisfaction of different user populations). These insight-gaining type of simulations can answer questions like: What has the greatest influence on X? or How will X and Y interact?.
@@
Second, the study of user simulation was also motivated by the need for using reproducible experiments to evaluate an interactive information access system. While it is always desirable to use real users for evaluating an interactive system, it is hard to conduct reproducible experiments to fairly compare two systems or evaluate the impact of any component in a system. It is especially difficult to control for the fact that once a user interacts with a system to perform a task, the user’s cognitive state changes. The next time the same user might not interact the same way with that system, or with any other system, when performing the same or related tasks. It is also costly to use many real users to conduct large scale experiments. Building user simulators that can realistically simulate real users offers an attractive solution to this problem, since a simulated user can be “reused” again and again in a controlled manner to ensure the reproducibility of experiments without any bias. Moreover, when a simulated user has a meaningful user parameter, it would also allow for varying the parameter to simulate user variations. Indeed, the Cranfield evaluation methodology developed in the IR community in 1960s [Sanderson:2010:FnTIR] was based on such a user simulation methodology, only that the assumed user model was not explicitly articulated and was overly simplified. Later research attempted to adopt more realistic user models and define additional evaluation measures based on more accurate user simulation. Simulations with the explicit focus of evaluating systems enable exploration of changes in system performance under different conditions or behavioural patterns (“what-if” analyses).
@@
Finally, research on user simulation may also be motivated by the need to learn from user feedback to optimize a ranking or recommendation algorithm. For example, an effective ranking/recommendation algorithm would essentially attempt to accurately estimate the likelihood that a user would click on an information item. If the estimated clicking probability is high, the item should be ranked high on the result/recommendation list. This connection means that optimizing an algorithm and accurately modeling a user’s behaviour are in some sense the two sides of the same coin. As such, the models developed for optimizing ranking or recommendation can often be viewed as models for modeling a user’s clicking behaviour. Research on user simulation from this perspective generally uses sophisticated machine learning algorithms to combine multiple factors (predictors) that can all potentially help improve the accuracy prediction, which is appropriate for the purpose of optimizing an algorithm in a system, since the system does have access to all the information. However, from the perspective of simulating a real user, many of those factors cannot be assumed to be available to a user when choosing an action to interact with a system. Therefore, a click model optimized for a system may not be an optimal model for simulating a real user, but many techniques developed for optimizing ranking may be adapted for modeling a user’s clicking behaviour.
@@ 3.2 User Simulation for Search Engines
So far, most existing work on user simulation has been done in the context of search engines. In this section, we provide a detailed overview of this body of work. In the next section, we will provide a brief overview of user simulation in the context of recommender systems and conversational assistants, which reflects more what the recent research trend is.
The classic problem in information retrieval, ad hoc document retrieval, is to return a ranked list of documents from a collection in decreasing order of their estimated relevance to a user-provided search query. Typical system-oriented evaluation assumes a persistent scanning of a long list of search results, where the user reads each document before making a judgment about its relevance. In practice, the ways in which users interact with text-based information retrieval systems when conducting searches are a lot more complex. First, results are presented on a search engine results page (SERP), with features designed to help people find information faster. One of those features is snippets, which provide a “preview” of each search result by including the title and URL of the website, and a textual excerpt from the document that highlights its relevance to the search query. A searcher typically assesses the result snippet before clicking on them to read the full document. The lower the rank of the snippet, the higher the probability that they would stop examining the result list; in practice, few users scan past the second results page [Markey:2007:JASIST]. It might also happen that the searcher abandons the SERP based on an initial impression, without inspecting any results in depth. As a matter of fact, users often need to reformulate their queries to find what they are looking for. Over the course of the search session, the user may learn by reading snippets and inspecting documents, which might shape their view of relevance. The searcher will also at some point make a decision to stop their interaction with the search engine. Stopping may occur for a variety of reasons, for example, the searcher has satisfied their information need, has been frustrated with the system, or has run out of time.
There is a rich body of research on aspects that can influence the behaviour and performance of searchers. Below, we present a selection studies around three main themes that have been extensively investigated in the past: formulating queries, examining search results, and modeling searcher effort.
@@ 3.2.1 Formulating Queries
Azzopardi:2009:SIGIR studies the relationship between user effort, in terms of query length, and utility, measured with standard retrieval metrics (e.g., average precision). Under the Principle of Least Effort [Zipf:1949:Book], it is assumed that the user aims to minimize their cost/effort while maximizing their expected utility. An economic analysis of the productivity of querying is carried out using a set of simulated query-document pairs, which are produced using a generative probabilistic model (cf. Section 6.2). The empirical results show that queries of length between two and five are the most effective—this closely matches the lengths of queries observed from real users.
@@ 3.2.2 Examining Search Results
When inspecting the returned result list, users need to make a decision based on the result snippet as to whether they are interested in viewing the full document. If the snippet is deemed irrelevant, a relevant document may never get examined. Turpin:2009:SIGIR investigate the effects of including the snippet examination stage in the retrieval process and analyze how that might alter the relative ordering of different retrieval systems. It is observed that unless snippets are extremely accurate in how they reflect the content of the underlying document, they have a significant impact on system ordering. Turpin:2009:SIGIR therefore argue that snippets should be judged explicitly, in addition to documents, as part of the batch evaluation paradigm in order to model reality.
@@
The scanning strategy defines how the searcher interacts with the result list. A commonly assumed behaviour is persistent scanning, which means that the user examines all results in the specified range (rank position $k$), from top to down, in sequential order [Baskaya:2013:CIKM]. Often, scanning behaviour is considered from the opposite perspective of stopping behaviour: determining when the user would terminate the search [Kraft:1979:IPM, Maxwell:2019:PhDThesis]. Persistent scanning corresponds to stopping at a fixed depth, regardless of the relevance of results. This stopping strategy is encoded in several IR evaluation measures, for example, Precision@k and Normalized Discounted Cumulative Gain (NDCG). Clearly, fixed depth stopping is a simplified model of user behaviour. Intuitively, the user would decide when to stop scanning results based on their relevance. Maxwell:2015:CIKM explore how the performance of ad hoc document retrieval changes under different stopping strategies. For example, the searcher would stop once they have observed a given number of non-relevant snippets (total non-relevant) or a given number of non-relevant snippets in a row (contiguous non-relevant); cf. Section 6.6. A comparison of a total of six different stopping strategies in terms of mean Cumulative Gain reveals that the aforementioned two strategies perform best [Maxwell:2015:CIKM]. However, somewhat surprisingly, the baseline stopping strategy (fixed depth, with a well-chosen depth threshold) performs remarkably well, and has not been significantly outperformed by more advanced methods.

### v2

@@ 2.4 User Modeling
Conceptually, user modeling is needed in any system that attempts to serve a user, especially if it is desirable to make the service personalized to suit the needs of each individual user or optimize its interactions with an individual user. As interactive AI applications become increasingly feasible, user modeling is also attracting much attention with new research results regularly reported in multiple conferences such as ACM UMAP, [^2] https://www.um.org/conferences ACM IUI, [^3] https://iui.acm.org/ ACM CHI, [^4] https://chi.acm.org/chi-series/ and the IR-focused CHIIR. [^5] https://chiir.org/ From the perspective of user modeling, user simulation can be regarded as developing a complete and operational user model, which makes work on user modeling highly relevant to the discussion of user simulation.
@@
Within the broader field of user modeling, the most relevant lines of work for include the following: (1) Search tasks and intent, including, e.g., typology of search tasks [Marchionini:2006:ACM], taxonomy of Web search [Broder:2002:SIGIRForum], and taxonomy of e-commerce queries [Sondhi:2018:SIGIR]. These studies provide a basis for studying simulation of different kinds of search tasks and intent. (2) Information seeking models, including, e.g., exploratory information seeking models [Ellis:1989:JD, Meho:2003:JASIST]. Such models provide guidance on high-level workflow of a simulated user when performing an information seeking task. (3) Cognitive models of users, which attempt to model the knowledge state of a user; for example, the ASK theory [Belkin:1982:JD] and Ingwersen:1996:JD’s cognitive IR model [Ingwersen:1996:JD]. (4) Contextual user modeling[Ingwersen:2005:book], which takes a broader view of the context that may affect a user’s information seeking behaviour. (5) Click models[Chuklin:2015:Book]that can be directly used in a user simulator to simulate a user’s clicking behaviour. (6) Economic IR models[Azzopardi:2011:SIGIR, Azzopardi:2014:SIGIR], which provide guidance on how users make decisions during an interactive retrieval process and may also be directly used in a user simulator. (7) Choice modeling[Hazrati:2024:UMUAI], which leverage interpretable user simulators to simulate and analyze the overall behaviour of a community of users when interacting with a recommender system over a long time period.
As user simulation is essentially to define a computational model for a user, the discussion of user modeling will appear throughout the book. However, as a distinct research topic, user modeling plays an increasingly crucial role as we strive to optimize the collaboration between AI systems and human users. This is because the objective function to be optimized by an AI system in assistive applications must formally describe the user with whom the AI system collaborates. So far, however, the user simulation techniques only utilize a limited portion of research results on user modeling, leaving much room for future work on incorporating more user modeling research results into user simulation.
User modeling has also often been used as a way to study dynamic behaviour of a whole community of users interacting with a system. For example, the overall impact of a recommender system on a community of users under certain assumptions about user behaviour has been studied in some work [Yao:2023:ICML, Hazrati:2024:UMUAI]. In such studies, interpretable (often simplified) user models are often used to enable the simulation and analysis of how a system interacts with a community of users over time. This can be viewed as modeling the collective behaviour of a user community within a social environment, based on the individual models of each user. Our book is focused on discussing simulation of individual users in the context of their interactions with an interactive information access system.
@@
In this chapter, we provide a high-level broad overview of research on user simulation with a discussion of the different motivations and applications of how user simulation has been employed in the past. Specifically, we identify three different purposes of studying user simulation: (1) analysis of user behaviour in specific applications, (2) automatic evaluation of systems, (3) synthetic data generation for training models (e.g., using reinforcement learning). While our main interest in this book lies in (2), it is worth pointing out that the first two categories share a lot of similarities in terms of modeling user behaviour in an interpretable way; in contrast, the interpretability aspect may be less critical for the third category of application. Moreover, as the research in the first two categories has historically been done mostly in the context of search engine applications, our discussion in the rest of this chapter will also be mostly about search engines, though we will also include a brief discussion of recommender systems and conversational assistants, which have attracted much attention recently. Specifically, in the rest of the chapter, we will first discuss the different motivations for studying user simulation, roughly corresponding to different benefits of user simulation. We then provide an overview of the research done so far in the context of three representative information access systems, i.e., search engines, recommender systems, and conversational assistants. We end the chapter with a discussion of the different applications of user simulation, including particularly the use of simulation for evaluation, which is the main focus of this book.
@@ 3.1 Motivations for User Simulation
Analysis of user behaviour. The study of user simulation has often been done in the context of studying user behaviour, where user simulation can provide multiple benefits, such as formally defining a hypothesis about a user’s behaviour, modeling variations of user factors and their influences on user behaviour, and counterfactual analysis of impact of system components on user behaviour. In general, studying how various factors influence user interaction is important for both understanding users and optimizing the design of user interfaces and algorithms to serve users. For this reason, the study of user behaviour in interactive systems has been the subject of decades of user-oriented IR research (cf. Section 2.4). However, a major challenge in user studies is that the experimental conditions can introduce variables that are difficult to control (e.g., users’ background, domain knowledge, or perseverance), but they can affect how users interact with an information access system. Such uncontrolled variables can be explicitly modeled in user simulations to enable more accurate modeling of a user’s real behaviour.
@@
Moreover, simulation also allows for examining the effects of changes made to various system components on anticipated user behaviour. It facilitates a more comprehensive assessment than what a live environment would typically allow (e.g., estimating long-term engagement/satisfaction of different user populations that can be controlled or varied in a counterfactual manner). These insight-gaining type of simulations can answer questions like: What has the greatest influence on X? or How will X and Y interact?.
@@
Evaluation of interactive systems. The study of user simulation was also motivated by the need for using reproducible experiments to evaluate an interactive information access system. While it is always desirable to use real users for evaluating an interactive system, it is hard to conduct reproducible experiments to fairly compare two systems or evaluate the impact of any component in a system. It is especially difficult to control for the fact that once a user interacts with a system to perform a task, the user’s cognitive state changes. The next time the same user might not interact the same way with that system, or with any other system, when performing the same or related tasks. It is also costly to use many real users to conduct large scale experiments. Building user simulators that can realistically simulate real users offers an attractive solution to this problem, since a simulated user can be “reused” again and again in a controlled manner to ensure the reproducibility of experiments without any bias. Moreover, when a simulated user has a meaningful user parameter, it would also allow for varying the parameter to simulate user variations. Simulations with the explicit focus of evaluating systems enable exploration of changes in system performance under different conditions or behavioural patterns (“what-if” analyses).
@@
Implicit feedback (for machine learning). Research on user simulation may also be motivated by the need to learn from user feedback to optimize a ranking or recommendation algorithm. For example, an effective ranking/recommendation algorithm would essentially attempt to accurately estimate the likelihood that a user would click on an information item. If the estimated clicking probability is high, the item should be ranked high on the result/recommendation list. This connection means that optimizing an algorithm and accurately modeling a user’s behaviour are in some sense the two sides of the same coin. As such, the models developed for optimizing ranking or recommendation can often be viewed as models for modeling a user’s clicking behaviour. Research on user simulation from this perspective generally uses sophisticated machine learning algorithms to combine multiple factors (predictors) that can all potentially help improve the accuracy prediction, which is appropriate for the purpose of optimizing an algorithm in a system, since the system does have access to all the information. However, from the perspective of simulating a real user, many of those factors cannot be assumed to be available to a user when choosing an action to interact with a system. Therefore, a click model optimized for a system may not be an optimal model for simulating a real user, but many techniques developed for optimizing ranking may be adapted for modeling a user’s clicking behaviour.
Privacy protection. Study of user simulation may also address confidentiality concerns. Specifically, simulation enables us to experiment with, and evaluate, systems where the data and/or interactions are sensitive or confidential [Hawking:2020:Book]. There are many examples of applications where privacy needs to be protected, including, e.g., medical record search, email search, and enterprise search. In particular, simulation of Electronic Medical Records (EMRs) [Moniz:2009:JMIR] has been suggested and used in practice (see, e.g., MDClone [^1] https://www.mdclone.com/). If we have faith that a simulation accurately represents the salient aspects of some confidential setting, we can use that simulation for evaluation without ever seeing real data. Of course the accuracy of simulation still needs to be validated using the real data, but that validation can be performed by people who have access to the data.
User interface design. Finally, user simulation can also benefit user interface (UI) design as simulations would enable experimentation with changes that we do not want real users to see (yet), such as changes to the user interface of a running system. It is hard to make large-scale UI changes, since people are generally reluctant to learn new tools or even new interaction patterns, and the cost of increasing the cognitive burden on the real users due to change of an interface may be as high as encouraging the user to switch to a different system. Interface designers might also be reluctant to expose a new idea prematurely to real people, which may be another reason why we do not want real users to test a new interface feature. Simulation could potentially help address these challenges. For example, a carefully designed simulation may help assess the expected utility of a new interface, especially with the use of interpretable user simulators [Azzopardi:2011:SIGIR, Chi:2003:CHI, Fu:2007:HCI].
3.2 Applications of User Simulation
3.2.1 User Behaviour Analysis
Since a user simulator must be mathematically well defined, it serves as a formally specified hypothesis about a user’s behaviour. The benefit of a formally specified hypothesis is that the hypothesis can then be tested rigorously to examine its validity. This can be done by (1) designing focused user studies to examine to what extent the generated user behaviour of a formal user simulator match those of a real user or (2) fitting a formal user model to the observed real user data such as a search log to obtain estimates of the interpretable parameters in the formal model. While the user behaviour that can be studied in this way is limited to what can be modeled formally, this kind of research of user behaviour, which may be called “computational user modeling,” should presumably be our long-term goal in fully and precisely understanding a user. In this sense, advancement of user simulation generally should lead to improved understanding of users.
It is also interesting to note that the emergence of large language models (LLMs) opens up new opportunities of creating user simulators that can exhibit human-like behaviours especially in natural language conversations [Aher:2023:ICML]. While such LLMs are not interpretable, they have been shown to replicate some population behaviour based on some Turing Experiments [Aher:2023:ICML]. There is great potential to leverage LLMs for user simulation (see Chapter 9) with applications beyond evaluation (e.g., simulating complex social systems [Gao:2024:WWW]).
3.2.2 Evaluation of Interactive Systems
Second, virtually all the existing test collections (for both IR and recommender systems) can be regarded as an evaluation method based on some kind of simplified user simulators (see Chapter 4 for more discussion of this). A major limitation of those existing collections is that they cannot be used to evaluate an interactive system. Given the research progress already made in user simulation, we have the opportunity to address this limitation by systematically converting existing test collections into new collections with a wide range of user simulators trained based on the existing user data, thus enabling automatic evaluation of interactive search and recommender systems (see Section 10.1).
Finally, the application of user simulation for evaluation is closely related to its application to user studies in that in both cases interpretable user models are required in order for the applications to be successful. The requirement of interpretability in the case of user studies is obvious, but it is also required for evaluation, otherwise the evaluation results would not be so meaningful. With intrepretability, user simulation for evaluation would not only be meaningful but also be powerful in generating simulated users with different characteristics to allow for evaluation with different kinds of users. The integration of research of user studies and the development of formal interpretable user models is therefore highly desirable and may reinforce each other in accelerating research in both areas. Indeed, we envision a healthy ecosystem where user studies would lead to more realistic user simulators, which further enable the development of more effective interactive systems (e.g., with new functions and new interface features), leading to new opportunities of studying user behaviour in a new system environment (with new functions or new interfaces), and new generations of user simulators being developed.
3.2.3 Synthetic Data Generation
In addition to analysis of user behaviour and evaluation of systems, user simulation also has applications in synthetic data generation to create/augment training data for supervised machine learning, which tends to require large quantities of labeled examples of training data. Because manually annotating a data set is labor-intensive and costly, using user simulation to generate synthetic training data is appealing since it does not involve manual effort. Compared with other approaches to synthetic data generation, user simulation enables the generation of more interpretable synthetic data when the parameters used in the simulators are interpretable. Due to the scarcity of interaction data available (e.g., because of issues such as the privacy), it is especially beneficial to leverage user simulation to generate interaction data, which can then be used for training reinforcement learning algorithms for learning an effective interaction policy. Work on generation of synthetic data includes, but not limited to (1) modeling clicks and interactions between search results [Zhang:2019:CIKM], (2) training spoken dialogue systems [Schatzmann:2006:KER, Schatzmann:2007:NAACL, Young:2010:CSL], (3) generating synthetic conversations by leveraging large pre-trained language models [Mohapatra:2021:EMNLP, Kim:2021:ACL, Wan:2022:EMNLP], (4) generating a population of synthetic user profiles based on real user interaction data with items (simulation of tourists) [Merinov:2023:IIR], (5) removing bias from RL-based training of recommender systems [Huang:2020:RecSys], and (6) deep reinforcement learning for IR [Zhang:2021:SIGIR] and RecSys [Chen:2021:arXiv].
@@ 3.3 User Simulation for Search Engines
So far, most existing work on user simulation has been done in the context of search engines. In this section, we provide a detailed overview of this body of work. In the next two sections, we will provide a brief overview of user simulation in the context of recommender systems and conversational assistants, respectively, which reflects more what the recent research trend is.
The classic problem in information retrieval, ad hoc document retrieval, is to return a ranked list of documents from a collection in decreasing order of their estimated relevance to a user-provided search query. Typical system-oriented evaluation assumes a persistent scanning of a long list of search results, where the user reads each document before making a judgment about its relevance. In practice, the ways in which users interact with text-based information retrieval systems when conducting searches are a lot more complex. First, results are presented on a search engine results page (SERP), with features designed to help people find information faster. One of those features is snippets, which provide a “preview” of each search result by including the title and URL of the website, and a textual excerpt from the document that highlights its relevance to the search query. A searcher typically assesses the result snippet before clicking on them to read the full document. When a user sequentially browses a document list, the probability of stopping at a lower position is generally higher than stopping at a higher position; in practice, few users scan past the second results page [Markey:2007:JASIST]. It might also happen that the searcher abandons the SERP based on an initial impression, without inspecting any results in depth. As a matter of fact, users often need to reformulate their queries to find what they are looking for. Over the course of the search session, the user may learn by reading snippets and inspecting documents, which might shape their view of relevance. The searcher will also at some point make a decision to stop their interaction with the search engine. Stopping may occur for a variety of reasons, for example, the searcher has satisfied their information need, has been frustrated with the system, or has run out of time.
There is a rich body of research on aspects that can influence the behaviour and performance of searchers. Below, we present a selection of studies around three main themes that have been extensively investigated in the past: formulating queries, examining search results, and modeling searcher effort.
@@ 3.3.1 Formulating Queries
Azzopardi:2009:SIGIR studies the relationship between user effort, in terms of query length, and utility, measured with standard retrieval metrics (e.g., average precision). Under the Principle of Least Effort [Zipf:1949:Book], it is assumed that the user aims to minimize their cost/effort while maximizing their expected utility. [^2] Labhishetty:2022:ICTIR propose and study a unified optimization framework for modeling both initial query formulation and subsequent reformulations based on this principle. An economic analysis of the productivity of querying is carried out using a set of simulated query-document pairs, which are produced using a generative probabilistic model (cf. Section 6.2). The empirical results show that queries of length between two and five are the most effective—this closely matches the lengths of queries observed from real users.
@@ 3.3.2 Examining Search Results
When inspecting the returned result list, users need to make a decision based on the result snippet as to whether they are interested in viewing the full document. If the snippet is deemed irrelevant, a relevant document may never get examined. Turpin:2009:SIGIR investigate the effects of including the snippet examination stage in the retrieval process and analyze how that might alter the relative ordering of different retrieval systems. It is observed that unless snippets are extremely accurate in how they reflect the content of the underlying document, they have a significant impact on system ordering. Turpin:2009:SIGIR therefore argue that snippets should be judged explicitly, in addition to documents, as part of the batch evaluation paradigm in order to model reality. However, there has been little work on actually modeling a user’s interaction with snippets presumably due to the lack of such data sets.
The scanning strategy, also called a Browsing Model [Carterette:2011:SIGIR, Moffat:2013:CIKM, Moffat:2022:SIGIR], defines how the searcher interacts with the result list. Since the perceived utility of any search results is affected directly by how a user interacts with the results, it is not surprising that the browsing model has played an important role in defining evaluation metrics. Early work on evaluation measures has not explicitly specified a browsing model, making it unclear what assumptions have been made about users exactly. In the work  [Carterette:2011:SIGIR], the authors suggested a framework for examining an IR measure that consists of three components, including a browsing model, a model of document utility, and a model of accumulation of utility, allowing for systamtic comparison of different measures. In a series of work by Moffat and coauthors [Moffat:2008:TOIS, Moffat:2013:CIKM, Moffat:2017:TOIS, Moffat:2022:SIGIR], the browsing model is formally modeled with a probabilistic model, which enabled to derive a discounted co-efficient based on a browsing model when aggregating relevance information gain over search results. Specifically, with sequential browsing, the probability for a user to visit a result down on the list would generally be lower than the probability of visiting one ranked higher, leading to a natural discount of contribution of relevant information by a lowly ranked document as compared with a highly ranked one. The proposed frameworks, including the initial framework C/W/L [Moffat:2013:CIKM] and the extended one, C/W/L/A [Moffat:2022:SIGIR] provide a principled way to define evaluation measures based on an interpretable user model, including particularly an explicitly specified browsing model.
@@
In terms of specific browsing model, a commonly assumed scanning behaviour is persistent scanning, which means that the user examines all results in the specified range (rank position $k$), from top to down, in sequential order [Baskaya:2013:CIKM]. Often, scanning behaviour is considered from the opposite perspective of stopping behaviour: determining when the user would terminate the search [Kraft:1979:IPM, Maxwell:2019:PhDThesis]. Persistent scanning corresponds to stopping at a fixed depth, regardless of the relevance of results. This stopping strategy is encoded in several IR evaluation measures, for example, Precision@k and Normalized Discounted Cumulative Gain (NDCG). Clearly, fixed depth stopping is a simplified model of user behaviour. Intuitively, the user would decide when to stop scanning results based on their relevance. Maxwell:2015:CIKM explore how the performance of ad hoc document retrieval changes under different stopping strategies. For example, the searcher would stop once they have observed a given number of non-relevant snippets (total non-relevant) or a given number of non-relevant snippets in a row (contiguous non-relevant); cf. Section 6.6. A comparison of a total of six different stopping strategies in terms of mean Cumulative Gain reveals that the aforementioned two strategies perform best [Maxwell:2015:CIKM]. However, somewhat surprisingly, the baseline stopping strategy (fixed depth, with a well-chosen depth threshold) performs remarkably well, and has not been significantly outperformed by more advanced methods.